What is the price earning ratio for a hotel chain with 9 full service properties? The chain’s total revenue for 2008 was $79,238,000; the departmental expenses were 33,343,000; undistributed operating expenses were $17,345,000; rent, property taxes & insurance totaled $3,040,000; interest & depreciation expenses were $4,823,000; income tax was $6,206,000. The hotel’s had common shares outstanding at the beginning of the year 9,000,000 shares and at the end of the year 12,000,000 shares. The hotel’s average market price over the year was $23.12 per share.

Round your answer to two decimal points.

P/E ratio = ___
Answer #1

P/E ratio = MPS/EPS

MPS = $23.12

EPS = Income attributable to equity/common stock

(79,238,000 - 33,343,000 - 17,345,000 - 3,040,000 - 4,823,000 - 6,206,000)/(9,000,000 + 12,000,000)/2

= 14,481,000/10,500,000 = 1.38

P/E ratio = 23.12/1.38 = 16.75